OpenStudy (diyadiya):

Let f:R->R be a function defined by f(x) \[f(x)=- \frac{|x|^3+|x|}{1+x^2} \] then the graph lies in the A.1st & 2nd Quadrant B.1st & 3rd Quadrant C.2nd and 3rd Quadrant D.3rd & 4th Quadrant

OpenStudy (zarkon):

looks like f is always negative...

OpenStudy (zarkon):

or zero

OpenStudy (diyadiya):

so 3rd and fourth?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

since both numerator and denominator are positive, and there is a big fat minus sign out front

OpenStudy (diyadiya):

What about 2nd and 3rd?

OpenStudy (zarkon):

the graph is never in the 2nd quadrant

OpenStudy (zarkon):

the domain for this function is all reals...but the y values are zero or negative..thus just the 3rd and 4th quads

OpenStudy (zarkon):

to make you notation nicer you can use \[f:\mathbb{R}\to\mathbb{R}\] f:\mathbb{R}\to\mathbb{R} :)

OpenStudy (diyadiya):

Ok Thank you :D